Consider the reaction: 2KClO3 → 2KCl + 3O2
A sample of 5.45 L of oxygen is collected over water at a total pressure of
735.5 torr at 25°C. How many grams of oxygen have been collected? (at
25°C, the vapour pressure of water is 23.8 torr)
1
Expert's answer
2020-11-26T12:20:44-0500

According to the Dalton's Law of Partial Pressures, the total pressure in the container must be the sum of the pressures of the gas one collected and the water vapour:

"P_T = P_w + P_{O_2}" .

Therefore, the pressure of oxygen is:

"P_{O_2} = P_T-P_w = 735.5-23.8 = 711.7" torr.

Using the ideal gas law (ideal gas constant "R = 62.3637" L torr mol-1 K-1), the number of the moles of oxygen is:

"n = \frac{pV}{RT} = \frac{711.7\cdot5.45}{62.3637\cdot(273.15+25)} = 0.2086" mol.

The mass of oxygen collected is then the product of its number of the moles and its molar mass "M = 32.00" g/mol:

"m = nM = 0.2086\cdot32.00 = 6.675" g, or 6.68 g.

Answer: 6.68 g of oxygen have been collected.